Playgrounds are sold for p dollars each. The monthly revenue rom the sale of the model is given by R = 0.08p^2 + 300p. Revenue is the product o the price "p" and the demand (quantity sold). What is an expression for the monthly demand? What is the monthly emand for this playground when the price is $3000.
Found 2 solutions by stanbon, funmath:
Answer by stanbon(75887)

Playgrounds are sold for p dollars each. The monthly revenue rom the sale of the model is given by R = 0.08p^2 + 300p. Revenue is the product o the price "p" and the demand (quantity sold). What is an expression for the monthly demand? What is the monthly demand for this playground when the price is $3000.
R=p(0.08p+300)
Demand=0.08p+300
D(3000)=0.08(3000)+300
=540

Answer by funmath(2933)

Playgrounds are sold for p dollars each. The monthly revenue rom the sale of the model is given by R = 0.08p^2 + 300p. Revenue is the product o the price "p" and the demand (quantity sold). What is an expression for the monthly demand? What is the monthly emand for this playground when the price is $3000.
R=p*(demand)
R=0.08p^2+300p Factor out the p
R=p*(0.08p+300)
demand=0.08p+300 <---first answer
:
demand=0.08(3000)+300
demand=240+300
demand=540 <---second answer
:
Happy Calculating!!!